# Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataξ,γ          Muhammad Abdul-Mageedλ,ξ,γ

### v1

@@ Abstract
Enhancing the linguistic capabilities of Large Language Models (LLMs) to include low-resource languages is a critical research area. Current research directions predominantly rely on synthetic data generated by translating English corpora, which, while demonstrating promising linguistic understanding and translation abilities, often results in models aligned with source language culture. These models frequently fail to represent the cultural heritage and values of local communities. This work proposes a methodology to create both synthetic and retrieval-based pre-training data tailored to a specific community, considering its (i) language, (ii) cultural heritage, and (iii) cultural values. We demonstrate our methodology using Egyptian and Moroccan dialects as testbeds, chosen for their linguistic and cultural richness and current underrepresentation in LLMs. As a proof-of-concept, we develop NileChat, a 3B parameter LLM adapted for Egyptian and Moroccan communities, incorporating their language, cultural heritage, and values. Our results on various understanding, translation, and cultural and values alignment benchmarks show that NileChat outperforms existing Arabic-aware LLMs of similar size and performs on par with larger models. We share our methods, data, and models with the community to promote the inclusion and coverage of more diverse communities in LLM development. [^1] https://github.com/UBC-NLP/nilechat.
@@
Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataξ,γ          Muhammad Abdul-Mageedλ,ξ,γ λThe University of British Columbia           δUM6P           ψPSU ξMBZUAI           γInvertible AI {abdellah.elmekki,muhammad.mageed}@ubc.ca
{strip}[Uncaptioned image]
Figure 1: Our proposed framework for data augmentation tailored to low-resource local communities. (a) Workflow for generating educational data in the target language through machine translation. (b) Workflow for generating diverse text genres in the target language by simulating scenarios that incorporate local cultural contexts and persona descriptions. (c) Retrieval process for augmenting local cultural knowledge by parsing web pages representing specific local cultural concepts through web search.
@@ 1 Introduction
This paper addresses the critical need to adapt multilingual LLMs to low-resource language communities by jointly considering their linguistic characteristics and cultural heritage & values. We propose a novel pipeline (illustrated in Figure 1) focused on data augmentation for continued pre-training. Our approach combines controlled synthetic data generation (Section 3.1.2) with retrieval (Section 3.1.3) methods. To address linguistic adaptation, we translate English pre-training data into the target local language focusing only on high-quality data from the educational domain (Section 3.1.1). Crucially, to imbue cultural relevance, we generate diverse texts reflecting specific cultural heritage concepts (e.g., food, celebrations, proverbs) using local persona descriptions (Section 3.1.2) reflecting the local cultural values. We demonstrate our method on the Moroccan and Egyptian Arabic dialects as low-resource testbeds. We further pre-train a multilingual LLM on a curated mix of real and synthetic data, evaluating its performance on tasks involving language understanding, translation, and alignment with cultural knowledge and values. Our findings show that the adapted model significantly outperforms baseline and existing models that are even bigger in size on most evaluation tasks.
@@ 3.1 Data Augmentation
The construction of LLMs that can serve a specific population fundamentally depends on the availability of representative data. Recognizing the acute scarcity of publicly available pre-training corpora for many low-resource languages, including EGY and MOR, we propose a novel data production method encapsulating the LHV dimensions of a given country-level population. As depicted in Figure 1, our approach leverages three complementary strategies intended to collectively capture the LHV dimensions: (a) machine translation (MT), (b) controlled-generation and (c) retrieval. We explain these next.
@@ 3.1.1 MT for Knowledge and Fluency
3.1.2 Controlled Generation for Cultural Heritage and Cultural Values
Linguistic fluency, while a foundational capability for LLMs, does not inherently guarantee their awareness of, or alignment with, the culture and values of a specific target community Naous et al. (2024). To bridge this gap, we employ controlled generation. For controlled generation, we use the teacher LLM to generate diverse texts in the target language. These texts are specifically designed to discuss local topics, which are identified from articles sourced from local news websites or the target country’s Wikipedia portal. Furthermore, the generated content is crafted to reflect distinct personas, each defined by a profile encompassing specific moral values, demographic characteristics, and socioeconomic attributes. Our approach integrates four key components to achieve this:
@@ 3.1.2 Controlled Generation for Cultural Heritage and Cultural Values
Core Cultural Heritage Concepts. We integrate key regional cultural elements, such as cuisine, landmarks, and celebrations, by extracting relevant articles from country-specific Wikipedia portals.
@@
To produce diverse text genres for pre-training, we combine data points from the four listed components into a unified prompt to guide the teacher LLM. This prompt instructs the LLM to generate varied text outputs in the target low-resource language, explicitly integrating the selected persona’s values, the specified cultural concepts, and provided linguistic cues. Specifically, we focus on generating the following genres: stories, personal essays, blog posts, reviews, and conversations. An example of this process is depicted in Figure 1 (b).
@@ 3.2.1 Continued Pre-Training
We generate pre-training data for EGY and MOR using the methods outlined in Section 3.1. Our approach involves three main components: (i) MT Data. We employ our teacher model to translate English educational content into both dialects. Specifically, we sample 5.5 million texts from the Fineweb-edu dataset Penedo et al. (2024) and translate them into EGY and MOR. (ii) Controlled Generation Data. We craft tailored prompts incorporating personas, local cultural contexts, dialectal glosses, expressions, and utterances to instruct the LLM in generating diverse genres of text. For persona descriptions, we generate 1,200 descriptions based on data from Egyptian and Moroccan participants in the WVS. For local news context, we leverage an in-house corpus comprising approximately 1.5 million Egyptian and 800,000 Moroccan news articles, originally published in MSA by local news websites. Additionally, we include 25,000 Egyptian and 49,000 Moroccan Wikipedia articles. For dialectal glosses, expressions, and utterances, we draw from publicly available resources on EGY and MOR proverbs and idiomatic expressions, each accompanied by English explanations. We further augment this with an in-house dataset of 600 dialectal utterances from Egyptian and Moroccan television shows paired with English translations, as well as 4,000 dialect-to-English word pairs for each dialect from the Gatitos dictionary Jones et al. (2023). (iii) Retrieval Data. For information retrieval, we query the Brave Search API [^2] https://brave.com/search/api/ using 6,500 cultural concepts from Morocco and 4,500 cultural concepts from Egypt. These concepts represent the ten cultural heritage categories in the set {food, clothes, landmarks, festivals & celebrations, geography, handicrafts, architecture, auna, flora, music}.
@@
The generated dataset comprises approximately 5.5 million educational articles for both EGY and MOR. Additionally, for EGY, it includes approximately 300,000 samples for each category of conversions, personal essays, blog posts, reviews, and stories. For MOR, there are approximately 150,000 samples for each of these same categories. These latter categories represent the LHV dimensions (§3.1). Table A.3 presents a sample of these texts.
@@
A filtering process using a repetitive n-gram filter removed 3.97% of the data. We also conducted a dialectness check on the generated data using ALDi Keleg et al. (2023). The average dialectness scores for the EGY and MOR educational articles are $0.45$ and $0.32$, respectively. In contrast, for the texts focused on cultural heritage and values, the average dialectness scores are higher, at $0.84$ for EGY and $0.72$ for MOR. We attribute the lower dialectness levels in the educational articles to the prevalence of scientific terms that often lack direct equivalents in EGY and MOR, and were therefore retained in MSA. We convert 1.5M and 0.5M from this generated data to Arabizi for both EGY and MOR, respectively. For retrieval, we collect 110,000 and 30,000 articles about cultural heritage to both EGY and MOR.
@@
We used a cluster of 4$\times$×A100 80GB GPUs for 1,096 hours to create our augmented pre-training dataset using the listed inputs.
@@
Rather than pretraining an LLM from scratch, we continue pretraining Qwen-2.5-3B Qwen et al. (2025) with our data. We select this model due to its competitive performance and good tokenizer compression ratio on MSA. We continue pretraining the full model (3.1B parameters) for one whole epoch, which took 750 hours on 4$\times$×A100 80GB GPUs. More details about the base model selection and the training are in Appendix B.1.
@@ 3.2.2 Supervised Fine-Tuning
To adapt our pre-trained model for instruction following, we perform supervised fine-tuning (SFT). Due to the scarcity of SFT datasets for EGY and MOR, we construct a comprehensive training set. This process involves several key steps: (i) translation of SmolTalk dataset Allal et al. (2025) into MOR, EGY, French, and MSA using the teacher LLM; (ii) generation of synthetic dialectal question-answer pairs using our the retrieved dataset of local Egyptian and Moroccan cultural heritage; [^3] This data is initially created in Arabic script, and a portion is subsequently converted to Arabizi. (iii) incorporation of the Darija-SFT-Mixture MOR dataset provided by Shang et al. (2025); and (iv) translation of TULU-V2-mix dataset Ivison et al. (2023) into EGY. Finally, (v) this consolidated SFT dataset is augmented by converting understanding and generation tasks from the training sets of the ORCA Elmadany et al. (2023) and Dolphin Nagoudi et al. (2023) benchmarks into instruction-response formats. The final composition of our instruction dataset is in Table A.2.
@@
For model fine-tuning, we follow recent approaches Ramé et al. (2024); Dang et al. (2024) that leverage model merging techniques to produce models effective across multiple languages or tailored for particular tasks. Specifically, we fully fine-tuned two separate variants of the base model—one specialized for MOR and the other for EGY—each trained on its respective dialectal data in both Arabic script and Arabizi (plus an amount of shared data between the two variants from the other languages; see B.2). We fine-tune each dialect-specific model for two epochs and employ weighted linear averaging Aakanksha et al. (2024) for merging, dubbing our merged model NileChat.
@@ 4.1 Evaluation Tasks
We evaluate understanding capabilities using MMLU Hendrycks et al. (2021), HellaSwag Zellers et al. (2019), and Belebele Bandarkar et al. (2024) benchmarks, each adapted to both EGY and MOR dialects. For MOR, we directly employ the MMLU and HellaSwag versions provided by Shang et al. (2025). For EGY, we follow the translation pipeline described in Shang et al. (2025), translating the English and MSA MMLU tasks and the English HellaSwag dataset into EGY using our teacher model. For the Belebele benchmark, we utilize the official Moroccan and Egyptian dialect sets. Evaluations are conducted in both zero-shot and 3-shot scenarios, using accuracy as our performance metric.
@@
To assess cultural knowledge specific to Morocco and Egypt, we utilize the publicly available test set from the Palm benchmark Alwajih et al. (2025), focusing on these two countries. We adopt an LLM-as-Judge methodology Zheng et al. (2023), employing Gemma-3-27b Team et al. (2025b) to rate the correctness of model-generated responses compared to ground-truth answers on a scale from 0 to 10. The final evaluation score is calculated as the average correctness across all responses.
@@ 4.2 Results and Discussion
Less than 7B Qwen3-1.7B 28.53 28.53 28.44 27.47 22.89 22.89 3.61 2.12
@@
More than 7B AceGPT-7B-chat 40.29 37.57 33.27 30.47 32.67 32.00 5.58 3.93
@@
As shown in Table 1, our approach significantly enhances cultural knowledge (Palm), enabling NileChat to achieve scores of 5.72 (EGY) and 5.86 (MOR), compared to baseline Qwen2.5-3B-instruct scores of 2.86 and 2.31, respectively. Among similarly sized models, ours achieves the highest performance on MOR and ranks second only to Gemma-3-4B for EGY. Although larger models such as Gemma-3-12B exhibit superior overall scores (EGY: 8.71, MOR: 7.09), NileChat notably surpasses AceGPT-7B and -13B on Moroccan cultural knowledge, despite their despite their claimed alignment with Arabic cultures. Additionally, it outperforms Atlas-chat-2B and -9B, models specifically fine-tuned for Moroccan dialects. These results support our claim that linguistic fluency alone—gained through supervised fine-tuning or pre-training on potentially biased, translated datasets—is insufficient for genuine cultural alignment with local communities.
@@
XX $\rightarrow$ $\rightarrow$ XX XX $\rightarrow$ $\rightarrow$ XX
@@
Less than 7B Qwen3-1.7B 14.75 10.89 19.51 15.47 11.41 4.36 15.63 6.32 12.29
@@
More than 7B AceGPT-7B-chat 18.02 11.33 21.11 17.46 14.73 4.95 20.10 7.47 14.40
@@
Table 2: Zero-shot translation performance (spBLEU) on the Flores and in-house datasets. XX $\rightarrow$ EGY and XX $\rightarrow$ MOR denote average over target languages EGY and MOR, respectively. Conversely, EGY → XX and MOR → XX indicate average over EGY and MOR as source languages. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column. Detailed results are in Table D.2.

### v3

# Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataγ          Muhammad Abdul-Mageedλ,γ

@@ Abstract
Enhancing the linguistic capabilities of Large Language Models (LLMs) to include low-resource languages is a critical research area. Current research directions predominantly rely on synthetic data generated by translating English corpora, which, while demonstrating promising linguistic understanding and translation abilities, often results in models aligned with source language culture. These models frequently fail to represent the cultural heritage and values of local communities. This work proposes a methodology to create both synthetic and retrieval-based pre-training data tailored to a specific community, considering its (i) language, (ii) cultural heritage, and (iii) cultural values. We demonstrate our methodology using Egyptian and Moroccan dialects as testbeds, chosen for their linguistic and cultural richness and current underrepresentation in LLMs. As a proof-of-concept, we develop NileChat, a 3B parameter Egyptian and Moroccan Arabic LLM adapted for Egyptian and Moroccan communities, incorporating their language, cultural heritage, and values. Our results on various understanding, translation, and cultural and values alignment benchmarks show that NileChat outperforms existing Arabic-aware LLMs of similar size and performs on par with larger models. This work addresses Arabic dialect in LLMs with a focus on cultural and values alignment via controlled synthetic data generation and retrieval-augmented pre-training for Moroccan Darija and Egyptian Arabic, including Arabizi variants, advancing Arabic NLP for low-resource communities. We share our methods, data, and models with the community to promote the inclusion and coverage of more diverse communities in cultural LLM development. [^1] https://github.com/UBC-NLP/nilechat.
@@
Abdellah El Mekkiλthanks: Work initiated during a stay at MBZUAI.          Houdaifa Atouδ          Omer Nacarψ Shady Shehataγ          Muhammad Abdul-Mageedλ,γ λThe University of British Columbia           δMohammed VI Polytechnic University ψTuwaiq Academy           γInvertible AI {abdellah.elmekki,muhammad.mageed}@ubc.ca
@@ 1 Introduction
This paper addresses the critical need to adapt multilingual LLMs to low-resource language communities by jointly considering their linguistic characteristics and cultural heritage & values. We propose a novel pipeline (illustrated in Figure LABEL:fig:synthetic_pre_training) focused on data augmentation for continued pre-training. Our approach combines controlled synthetic data generation (Section 3.1.2) with retrieval (Section 3.1.3) methods. To address linguistic adaptation, we translate English pre-training data into the target local language focusing only on high-quality data from the educational domain (Section 3.1.1). Crucially, to imbue cultural relevance, we generate diverse texts reflecting specific cultural heritage concepts (e.g., food, celebrations, proverbs) using local persona descriptions (Section 3.1.2) reflecting the local cultural values. We demonstrate our method on the Moroccan and Egyptian Arabic dialects as low-resource testbeds. We further pre-train a multilingual LLM on a curated mix of real and synthetic data, evaluating its performance on tasks involving language understanding, translation, and alignment with cultural knowledge and values. Our findings show that the adapted model significantly outperforms baseline and existing models that are even bigger in size on most evaluation tasks.
@@ 3.1 Data Augmentation
The construction of linguistically-rich and culturally-rich LLMs that can serve a specific population fundamentally depends on the availability of representative data. Recognizing the acute scarcity of publicly available pre-training corpora for many low-resource languages, including EGY and MOR, we propose a novel data production method encapsulating the LHV dimensions of a given country-level population. As depicted in Figure LABEL:fig:synthetic_pre_training, our approach leverages three complementary strategies intended to collectively capture the LHV dimensions: (a) machine translation (MT), (b) controlled synthetic data generation and (c) retrieval. We explain these next.
@@ 3.1.1 MT for Knowledge and Fluency
3.1.2 Controlled Synthetic Data Generation for Cultural Heritage and Cultural Values
Linguistic fluency, while a foundational capability for LLMs, does not inherently guarantee their awareness of, or alignment with, the culture and values of a specific target community Naous et al. (2024). To bridge this gap, we employ controlled synthetic data generation. For controlled generation, we use the teacher LLM to generate diverse texts in the target language. These texts are specifically designed to discuss local topics, which are identified from articles sourced from local news websites or the target country’s Wikipedia portal. Furthermore, the generated content is crafted to reflect distinct personas, each defined by a profile encompassing specific moral values, demographic characteristics, and socioeconomic attributes. Our approach integrates four key components to achieve this:
@@ 3.1.2 Controlled Synthetic Data Generation for Cultural Heritage and Cultural Values
Core Cultural Heritage Concepts. We integrate key local cultural elements, such as cuisine, landmarks, and celebrations, by extracting relevant articles from country-specific Wikipedia portals.
@@
To produce diverse text genres for pre-training, we combine data points from the four listed components into a unified prompt to guide the teacher LLM. This prompt instructs the LLM to generate varied text outputs in the target low-resource language, explicitly integrating the selected persona’s values, the specified cultural concepts, and provided linguistic cues. Specifically, we focus on generating the following genres: stories, personal essays, blog posts, reviews, and conversations. An example of this process is depicted in Figure LABEL:fig:synthetic_pre_training (b).
@@ 3.2.1 Continued Pre-Training
We generate pre-training data for EGY and MOR using the methods outlined in Section 3.1. Our approach involves three main components: (i) MT Data. We employ our teacher model to translate English educational content into both dialects. Specifically, we sample 5.5 million texts from the Fineweb-edu dataset Penedo et al. (2024) and translate them into EGY and MOR. (ii) Controlled Synthetic Generation Data. We craft tailored prompts incorporating personas, local cultural contexts, dialectal glosses, expressions, and utterances to instruct the LLM in generating diverse genres of text. For persona descriptions, we generate 1,200 descriptions based on data from Egyptian and Moroccan participants in the WVS. For local news context, we leverage an in-house corpus comprising approximately 1.5 million Egyptian and 800,000 Moroccan news articles, originally published in MSA by local news websites. Additionally, we include 25,000 Egyptian and 49,000 Moroccan Wikipedia articles. For dialectal glosses, expressions, and utterances, we draw from publicly available resources on EGY and MOR proverbs and idiomatic expressions, each accompanied by English explanations. We further augment this with an in-house dataset of 600 dialectal utterances from Egyptian and Moroccan television shows paired with English translations, as well as 4,000 dialect-to-English word pairs for each dialect from the Gatitos dictionary Jones et al. (2023). (iii) Retrieval Data. For information retrieval, we query the Brave Search API [^2] https://brave.com/search/api/ using 6,500 cultural concepts from Morocco and 4,500 cultural concepts from Egypt. These concepts represent the ten cultural heritage categories in the set {food, clothes, landmarks, festivals & celebrations, geography, handicrafts, architecture, fauna, flora, music}.
@@
The generated dataset comprises approximately 5.5 million educational articles for both EGY and MOR. Additionally, for EGY, it includes approximately 300,000 samples for each category of conversations, personal essays, blog posts, reviews, and stories. For MOR, there are approximately 150,000 samples for each of these same categories. These latter categories represent the LHV dimensions (§3.1). Table A.3 presents a sample of these texts.
@@
A filtering process using a repetitive n-gram filter removed 3.97% of the data. We also conducted a dialectness check on the generated data using ALDi Keleg et al. (2023). The average dialectness scores for the EGY and MOR educational articles are $0.45$ and $0.32$, respectively. In contrast, for the texts focused on cultural heritage and values, the average dialectness scores are higher, at $0.84$ for EGY and $0.72$ for MOR. We attribute the lower dialectness levels in the educational articles to the prevalence of scientific terms that often lack direct equivalents in EGY and MOR, and were therefore retained in MSA. We convert 1.5M EGY and 0.5M MOR samples from the generated data to Arabizi. For retrieval, we collect 110,000 and 30,000 articles about cultural heritage for both EGY and MOR.
@@
[…]
@@
[…]
@@ 3.2.2 Supervised Fine-Tuning
To adapt our pre-trained model for instruction following, we perform supervised fine-tuning (SFT).
@@
Due to the scarcity of SFT datasets for EGY and MOR, we construct a comprehensive training set. This process involves several key steps: (i) translation of SmolTalk dataset Allal et al. (2025) into MOR, EGY, French, and MSA using the teacher LLM; (ii) synthetic generation of dialectal question-answer pairs using our retrieved dataset of local Egyptian and Moroccan cultural heritage; [^3] This data is initially created in Arabic script, and a portion is subsequently converted to Arabizi. (iii) incorporation of the Darija-SFT-Mixture MOR dataset provided by Shang et al. (2025); and (iv) translation of TULU-V2-mix dataset Ivison et al. (2023) into EGY. Finally, (v) this consolidated SFT dataset is augmented by converting understanding and generation tasks from the training sets of the ORCA Elmadany et al. (2023) and Dolphin Nagoudi et al. (2023) benchmarks into instruction-response formats. The final composition of our instruction dataset is in Table A.2.
@@
For model SFT, we follow recent approaches Ramé et al. (2024); Dang et al. (2024) that leverage model merging techniques to produce models effective across multiple languages or tailored for particular tasks. Specifically, we fully fine-tuned two separate variants of the base model—one specialized for MOR and the other for EGY—each trained on its respective dialectal data in both Arabic script and Arabizi (plus an amount of shared data between the two variants from the other languages; see B.2). We fine-tune each dialect-specific model for two epochs and employ weighted linear averaging Aakanksha et al. (2024) for merging, dubbing our merged model NileChat.
@@ 4.1 Evaluation Tasks
We evaluate understanding capabilities using MMLU Hendrycks et al. (2021), HellaSwag Zellers et al. (2019), and Belebele Bandarkar et al. (2024) benchmarks, each adapted to both EGY and MOR dialects. For MOR, we directly employ the MMLU and HellaSwag versions provided by Shang et al. (2025). For EGY, we follow the translation pipeline described in Shang et al. (2025), translating the English and MSA MMLU tasks and the English HellaSwag dataset into EGY using our teacher model  [^4] We have publicly released the EgyMMLU and EgyHellaswag benchmarks for evaluation on the LM Evaluation Harness Framework at https://github.com/EleutherAI/lm-evaluation-harness/tree/main/lm_eval/tasks/egymmlu and https://github.com/EleutherAI/lm-evaluation-harness/tree/main/lm_eval/tasks/egyhellaswag.. A careful verification of the translation quality for the generated EGY MMLU and EGY HellaSwag shows that the average correctness is 3.85 on a scale from 1-5 and the average dialectness is approximately 4 on a scale from 1-5. Further details are provided in Appendix C.1. For the Belebele benchmark, we utilize the official Moroccan and Egyptian dialect sets. Evaluations are conducted in both zero-shot and 3-shot scenarios, using accuracy as our performance metric.
@@
To assess cultural knowledge specific to Morocco and Egypt, we utilize the publicly available test set from the Palm benchmark Alwajih et al. (2025), focusing on these two countries only. We adopt an LLM-as-Judge methodology Zheng et al. (2023), employing Gemma-3-27b Team et al. (2025b) to rate the correctness of model-generated responses compared to ground-truth answers on a scale from 0 to 10. The final evaluation score is calculated as the average correctness across all responses.
@@ 4.2 Results and Discussion
Less than 7B
Qwen3-1.7B 28.53 28.53 28.44 27.47 22.89 22.89 3.61 2.12
@@
More than 7B
AceGPT-7B-chat 40.29 37.57 33.27 30.47 32.67 32.00 5.58 3.93
@@
As shown in Table 1, our approach significantly enhances cultural knowledge (Palm), enabling NileChat to achieve scores of 5.72 (EGY) and 5.86 (MOR), compared to baseline Qwen2.5-3B-instruct scores of 2.86 and 2.31, respectively. Among similarly sized models, ours achieves the highest performance on MOR and ranks second only to Gemma-3-4B for EGY. Although larger models such as Gemma-3-12B exhibit superior overall scores (EGY: 8.71, MOR: 7.09), NileChat notably surpasses AceGPT-7B and -13B on Moroccan cultural knowledge, despite their claimed alignment with Arabic cultures. Additionally, it outperforms Atlas-chat-2B and -9B, models specifically fine-tuned for Moroccan dialects. These results support our claim that linguistic fluency alone—gained through supervised fine-tuning or pre-training on potentially biased, translated datasets—is insufficient for genuine cultural alignment with local communities.
@@
$X$ $\rightarrow$ $\rightarrow$ $X$ $X$ $\rightarrow$ $\rightarrow$ $X$
@@
Less than 7B
Qwen3-1.7B 14.75 10.89 19.51 15.47 11.41 4.36 15.63 6.32 12.29
@@
More than 7B
AceGPT-7B-chat 18.02 11.33 21.11 17.46 14.73 4.95 20.10 7.47 14.40
@@
Table 2: Zero-shot translation performance (spBLEU) on the Flores and in-house datasets. $X$ $\rightarrow$ EGY and $X$ $\rightarrow$ MOR denote average over target languages EGY and MOR, respectively. Conversely, EGY → $X$ and MOR → $X$ indicate average over EGY and MOR as source languages. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column. Detailed results are in Table D.2.